# Scene Graph Generation in Large-Size VHR Satellite Imagery: A Large-Scale Dataset and A Context-Aware Approach

### v2

@@ 2.2.1 Scene Graph Generation Methods in Natural Imagery
Existing SGG methods for natural imagery have been dominated by the two-stage process consisting of OBD and relationship prediction between objects. Given the need for compatibility between OBD models and relationship prediction models within the SGG framework, the majority of existing SGG models, in alignment with the literature [43, 44], adopt Faster R-CNN as the standard for the OBD task. To generate high-quality scene graphs from images, a series of works have been explored in several directions, such as task-specific [36, 45, 46, 47, 48] and dataset-specific [43, 49, 50]. Notably, aggregating contextual information between objects has been shown to be effective for SGG. Xu et al. [36] first combined contextual information to refine the features of objects and relationships by using GRU [51] to pass messages between graphs. Based on the discovery that there are some inherent relationship patterns in subgraphs, Motifs [45] constructed a new global context computing mechanism. A series of context-based studies [52, 53, 54] have consecutively explored the messaging mechanisms of objects and relationships. Most recently, HetSGG [48] attempted to capture the relationships between objects in more detail by proposing a heterogeneous graph and proved its effectiveness. However, the above methods of exhaustive enumeration based on all pairs tend to result in excessive computation, which makes it difficult to run under common computational resources (e.g., 24 GB of memory in a single GPU) for large-size VHR SAI scenarios with a large number of objects and more object pairs.
@@ 2.2.2 Scene Graph Generation Methods in SAI
Therefore, it is necessary to develop a novel SGG method that can effectively select pairs containing rich knowledge as well as correctly predict relationship types of these pairs in large-size VHR SAI.
3 Details of the RSG Dataset
@@ 3.1.1 Image Collection
To meet the needs of practical applications, images in the RSG dataset are collected from Google Earth, with a spatial resolution ranging from 0.15m to 1m. Considering the value of cognitive understanding of large-size VHR SAI scenarios in practical applications closely related to human activities, such as transportation, energy, and life, we collected scenarios from more than 1,200 airports, ports, wind power stations, nuclear power stations, thermal power stations, construction sites, sports, service areas, toll stations, traffic bridges, and dams from all over the world. Unlike objects in natural imagery, objects in SAI are oriented in a variety of directions. Therefore, HBB-based annotation cannot provide accurate spatial information for oriented objects. To wrap objects more accurately and develop more suitable algorithms for oriented objects in SAI, all object instances in the RSG dataset are annotated with OBB. Samples of annotated instances in the RSG dataset can be seen in Fig. 2.
@@ 3.1.2 Annotation Criteria
For the relationship description, unlike the ”human-eye view” perspective of natural imagery, the ”birds-eye view” perspective in SAI makes the semantic relationships between objects absolute, so we discard the orientation words such as ”east”, ”south”, ”west”, and ”north”. Moreover, we have learned a lesson from the ambiguous labeling of SGG datasets in natural imagery and removed relationship descriptions whose semantics are obviously ambiguous. After careful screening, the semantic relationships are defined into 8 major categories: ”distance warning”, ”spatial topology”, ”functional description”, ”circuit layout”, ”movement status”, ”emission status”, ”construction status” and ”parking status”, which contain 58 subcategories as in Fig. 3(b).
In addition, Fig. 4 gives the interaction mapping between objects and relationships. There are eight colors in the figure, representing the eight major categories of relationships consistent with Fig. 3(b). It can be seen that almost every relationship category is associated with multiple object categories, which reflects the complex interactions between objects in large-size VHR SAI.
@@ 3.1.3 Annotation Management
RSG dataset has a standardized dataset construction pipeline: pre-annotation and rules refinement stage, large-scale detailed annotation stage, and quality-checking stage. In the pre-annotation stage, 6 experts in the SAI field are first invited to formulate annotation rules, and then the main authors pre-annotate the images with 11 types of scenarios, which are examined and evaluated by these experts, thereby forming a detailed object and relationship annotation document. During the large-scale annotation stage, the main authors provide comprehensive training to 9 professionals with rich interpretation experience in SAI, and each image is annotated by 2 professionals and checked by 1 professional at the same time. During the quality-checking stage, cross-sampling and revisions were performed by main authors.
@@ 3.2.1 Data Characteristics
Compared with the existing OBD dataset in SAI, the RSG dataset has unique properties on task diversity, category diversity, and scale diversity. Sparse high-value pairs, high intra-class variation and inter-class similarity are also important characteristics of the RSG dataset.
Figure 5: Examples of intra-class variations and inter-class similarities in relationship on the RSG dataset.
@@
Task diversity. RSG supports both the OBD task and SGG task in large-size VHR SAI. After processing the annotated images, it can further satisfy the various task requirements such as OBD based on HBB/OBB in large-size VHR SAI, SGG based on HBB/OBB in large-size VHR SAI, and so on.
Category diversity. RSG has various scene categories, various object categories, and various relationship categories. Specifically, it contains 11 categories of VHR SAI scenarios over the world, 48 categories of important objects, and 58 categories of high-value relationships, which will bring new opportunities for cognitive understanding of SAI.
Scale diversity. In the RSG dataset, there are many typical objects with extreme aspect ratios (e.g., runways, bridges, gravity dams), and there are large scale variations between objects of the same category (e.g., airplanes, ships) and different categories (e.g., ”boats and docks”, ”cars and car parkings”) that contain rich knowledge, which will bring great challenges to the OBD and SGG tasks in large-size VHR SAI.
Sparse high-value pairs. There are an enormous number of objects in a large-size SAI. If relationship prediction is performed for all object pairs, more than $N(N-1)$ triplets will appear for $N$ objects in the case of one-to-many relationships (e.g., in the RSG dataset where the number of objects in one image is up to 6,800, the possible triplets will be more than 40,000,000), which will cause memory overflow under common computational resources. In fact, not all object pairs are of interest to us, and only a few of them contain high-value relationships. Therefore, it will be a new challenge to select high-value pairs from the numerous pairs for the SGG task in large-size VHR SAI.
@@ 3.2.2 Data Statistics and Visualization
The statistical results of object and relationship annotations from the RSG dataset are shown in Fig. 3(a) and Fig. 3(b), respectively. It can be seen that there is an obvious long-tail effect for both object and relationship categories in the RSG dataset, which is also consistent with reality. The highest-frequency objects are common small objects such as ”cars”, ”airplanes”, ”tanks” and ”boats”. The highest-frequency relationships are mainly related to the parking states, such as ”parked alongside with”, ”parking in the different apron with”, ”parallelly parked on” and ”parking in the same apron with”.
@@ 4.1 Oriented Object Detection in Large-size VHR SAI
OBD in large-size SAI is an indispensable basis for SGG in large-size SAI, which directly affects the prediction accuracy of the triplets. Compared with small-size or low-resolution imagery, large-size VHR SAI can not only accommodate more complete large objects (e.g., runways, docks), but also greatly retain important details of small objects (such as the head and tail of a car, the head and tail of an airplane), which can help to mine more high-value knowledge in SAI. In processing large-size VHR SAI, mainstream deep learning-based OBD methods, whether supervised [56, 57, 58, 59] or weakly supervised [60, 61, 62, 63], adopt a cropping strategy [4, 64], which results in large objects, such as runways, being cut off. In addition to the cropping strategy, some OBD methods tackle the original large-size imagery with a fixed-window down-sampling strategy [65, 66, 67], resulting in the loss of massive detail information of the image. Recently, HBD-Net for single-class bridge detection in large-size VHR SAI is proposed [27], and the SDFF architecture in HBD-Net effectively solves the large scale variations in single-class bridge detection by fusing multi-scale contexts via the dynamic image pyramid (DIP) of the large-size image. However, this single-class OBD only considers the intra-class variations of the objects, which is not applicable in the multi-class OBD task with both intra-class and inter-class variations.
To address the aforementioned problem, this paper extends HBD-Net to the holistic multi-class OBD in large-size VHR SAI, named HOD-Net. In HOD-Net, the total loss of the HBB-based and OBB-based detectors in large-size VHR SAI is defined as follows.
$\mathcal{L}_{O}$ $=\sum_{m=1}^{\mathcal{M}}(\frac{1}{\Gamma^{m}}\!\sum_{i\in\Delta^{m}}\!\mathcal{L}_{i}^{cls}\!+\!\frac{1}{{\Gamma^{m}}^{+}}\sum_{j\in{\Delta^{m}}^{+}}\!w^{reg}_{j}\mathcal{L}_{j}^{reg}),$ (2)
$\mathcal{L}_{H}$ $=\sum_{m=1}^{\mathcal{M}}(\frac{1}{\Gamma^{m}}\!\sum_{i\in\Delta^{m}}\!\mathcal{L}_{i}^{cls}\!+\!\frac{1}{{\Gamma^{m}}^{+}}\!\sum_{j\in{\Delta^{m}}^{+}}\!\mathcal{L}_{j}^{reg}),$ (3)
@@ 4.2 Pair Proposal Generation via Adversarial Reconstruction
The training objective of $PED_{1}$ is to minimize the reconstruction error $\left\|X_{i}-X_{i}^{1}\right\|_{2}$ to learn the latent representation of the data, and the training objective of $PED_{2}$ is to reduce the reconstruction error $\left\|X_{i}-X_{i}^{2}\right\|_{2}$ to a minimum. As the iterations $n$ are constantly changing, the loss function is denoted as:
@@
4.3 Relationship Prediction with Context-Aware Messaging
@@ 4.3.1 Progressive Bi-context Augmentation Network
Figure 7: Illustration of the proposed PBA module. This illustration shows the details of messaging and fusion for one progressive context augmentation of entity-relationship.
@@
where $\sigma(\cdot)$ denotes the sigmoid activation function [71]. $\alpha^{ee},\alpha^{rs}$ and $\alpha^{ro}$ represent the importance of messaging from entity to entity, relationship to subject, and relationship to object, respectively. By learning to adjust $\alpha$ [70], each iteration leads to a change in attention and affects subsequent iters. $w^{ee},w^{rs}$ and $w^{ro}$ denote the weight matrix of the directed messaging from entity to entity, relationship to subject, and relationship to object, respectively. ${F}_{i}^{e}$ and ${F}_{i}^{r}$ are entity representation and relationship representation, respectively.
@@
Global-Local Feature Fusion. Introducing reliable global information helps to better recognize relationships, but more fine-grained discriminative features mainly come from local. Therefore, we adopt a global-local feature fusion strategy to capture reliable global information and maintain the local more discriminative features. Specifically, the context augmented feature of entity ${F}_{i+1}^{e}$ can be obtained by fusing the context-aware feature of entity $\widehat{F}_{i+1}^{e}$ from global and the entity feature ${F}_{0}^{e}$ from local. Similarly, the context augmentation feature of relationship ${F}_{i+1}^{r}$ can be obtained. After $L$ iters, the predicted relationship representations $Rel$ with context-aware augmentation are updated. It is worth noting that the predicted relationship representations in 4.3.1 are optimized and updated simultaneously with the prototype representations in 4.3.2.
@@ 4.3.2 Prototype-guided Relationship Learning
Prototype learning allows SGG models to focus on all categories equally during training, and is effective in alleviating the category imbalance problem of datasets [47]. However, previous linear classifier-based prediction methods often lack the ability to model data at different semantic levels and suffered from a lag in representational capabilities (e.g., $<$ship, away from, dock$>$ and $<$ship, approach, dock$>$ correspond to different relationship labels, but it is clear that $<$ship, away from, dock$>$ and $<$ship, approach, dock$>$ are closer than $<$truck, away from, gas station$>$ in the visual representation). We propose a prototype-guided relationship learning method that utilizes category labels as prompts to help generate learnable relationship prototypes, and prototype representations updated dynamically by instance-level constraints and prototype-level constraints to help the SGG model enhance the discriminability between prototypes.
@@
$\mathcal{L}_{IC}$ $=-log\frac{exp(<\overline{r},\overline{p}>/\tau)}{\textstyle\sum_{i=0}^{C}exp(<\overline{r},\overline{p}_{i}>/\tau)},$ (8)
$\mathcal{L}_{ID}$ $=max(0,q^{+}-q^{-}+\gamma_{1}),$ (9)
@@
In the training stage, the overall loss function $\mathcal{L}$ of our RPCM is denoted as:
@@
In the inference stage, the prototype type with the highest cosine similarity $s_{i}$ will be considered as the result of the relationship prediction:
@@
TABLE II: Baseline results (%) of HBB-based and OBB-based detectors on RSG test set. b_b denotes boarding_bridge, l_t denotes lattice_tower, s_l is ship_lock, and g_d represents gravity_dam. All experiments are based on the standard ‘1x’ (12 epochs) training schedule. means that Swin-L [72] is used and others indicates ResNet50 [73] as the backbone. Underline indicates the base detector for subsequent lines.
@@
TABLE III: Results (%) of different training strategies on RSG test set. b_b denotes boarding_bridge, l_t denotes lattice_tower, s_l is ship_lock, and g_d represents gravity_dam. All experiments are based on the standard ‘1x’ (12 epochs) training schedule. Swin-L [72] is adopted as the backbone.
@@
Figure 8: The visualization results of OBB and HBB detection on the RSG dataset using the HOD-Net and comparison OBD methods.

### v3

@@ 2.2.1 Scene Graph Generation Methods in Natural Imagery
Existing SGG methods for natural imagery have been dominated by the two-stage process consisting of OBD and relationship prediction between objects. Given the need for compatibility between OBD models and relationship prediction models within the SGG framework, the majority of existing SGG models, in alignment with the literature [41, 42], adopt Faster R-CNN as the standard for the OBD task. To generate high-quality scene graphs from images, a series of works have been explored in several directions, such as task-specific [34, 43, 44, 45, 46] and dataset-specific [41, 47, 48]. Notably, aggregating contextual information between objects has been shown to be effective for SGG. Xu et al. [34] first combined contextual information to refine the features of objects and relationships by using GRU to pass messages between graphs. Based on the discovery that there are some inherent relationship patterns in subgraphs, Motifs [43] constructed a new global context computing mechanism. A series of context-based studies [49, 50, 51] have consecutively explored the messaging mechanisms of objects and relationships. Most recently, HetSGG [46] attempted to capture the relationships between objects in more detail by proposing a heterogeneous graph and proved its effectiveness. However, the above methods of exhaustive enumeration based on all pairs tend to result in excessive computation, which makes it difficult to run under common computational resources (e.g., 24 GB of memory in a single GPU) for large-size VHR SAI with a large number of objects and more object pairs.
@@ 2.2.2 Scene Graph Generation Methods in SAI
3 Details of the STAR Dataset
@@ 3.1.1 Image Collection
To meet the needs of practical applications, images in the STAR dataset are collected from Google Earth, with a spatial resolution ranging from 0.15m to 1m. Considering the value of cognitive understanding of large-size VHR SAI scenarios in practical applications closely related to human activities, such as transportation, energy, and life, we collected scenarios from more than 1,200 airports, ports, wind power stations, nuclear power stations, thermal power stations, construction sites, sports, service areas, toll stations, traffic bridges, and dams from all over the world. Unlike objects in natural imagery, objects in SAI are oriented in a variety of directions. Therefore, HBB-based annotation cannot provide accurate spatial information for oriented objects. To wrap objects more accurately and develop more suitable algorithms for oriented objects in SAI, all object instances in the STAR dataset are annotated with OBB. Samples of annotated instances in the STAR dataset can be seen in Fig. 2.
@@ 3.1.2 Annotation Criteria
For the relationship description, unlike the “human-eye view” perspective of natural imagery, the “birds-eye view” perspective in SAI makes the semantic relationships between objects absolute, so we discard the orientation words such as “east”, “south”, “west”, and “north”. Moreover, we have learned a lesson from the ambiguous labeling of SGG datasets in natural imagery and removed relationship descriptions whose semantics are obviously ambiguous. After careful screening, the semantic relationships are defined into 8 major categories: “distance warning”, “spatial topology”, “functional description”, “circuit layout”, “movement status”, “emission status”, “construction status” and “parking status”, which contain 58 subcategories as in Fig. 3(b).
Fig. 4 gives the interaction mapping between objects and relationships. There are eight colors, representing the eight major categories of relationships consistent with Fig. 3(b). It can be seen that almost every relationship category is associated with multiple object categories, which reflects the complex interactions between objects in large-size VHR SAI.
@@ 3.1.3 Annotation Management
STAR dataset has a standardized dataset construction pipeline: pre-annotation and rules refinement stage, large-scale detailed annotation stage, and quality-checking stage. In the pre-annotation stage, 6 experts in the SAI field are first invited to formulate annotation rules, and then the main authors pre-annotate the images with 11 types of scenarios, which are examined and evaluated by these experts, thereby forming a detailed object and relationship annotation document. During the large-scale annotation stage, the main authors provide comprehensive training to 9 professionals with rich interpretation experience in SAI, and each image is annotated by 2 professionals and checked by 1 professional at the same time. During the quality-checking stage, cross-sampling and revisions were performed by main authors.
@@ 3.2.1 Data Characteristics
Compared with the existing OBD dataset in SAI, the STAR dataset has unique properties on task diversity, category diversity, and scale diversity. Sparse high-value pairs, high intra-class variation and inter-class similarity are also important characteristics of the STAR dataset.
Figure 5: Examples of intra-class variations and inter-class similarities in relationship on the STAR dataset.
@@
Task diversity. STAR supports both the OBD and SGG tasks in large-size VHR SAI. After processing the annotated images, it can further satisfy the various task requirements like OBD based on HBB/OBB in large-size VHR SAI, SGG based on HBB/OBB in large-size VHR SAI.
Category diversity. STAR has various scene categories, various object categories, and various relationship categories. Specifically, it contains 11 categories of VHR SAI scenarios over the world, 48 categories of important objects, and 58 categories of high-value relationships, which will bring new opportunities for cognitive understanding of SAI.
Scale diversity. In the STAR dataset, there are many typical objects with extreme aspect ratios (e.g., runways, bridges, gravity dams), and there are large scale variations between objects of the same category (e.g., airplanes, ships) and different categories (e.g., “boats and docks”, “cars and car parkings”) that contain rich knowledge, which will bring great challenges to the OBD and SGG tasks.
Sparse high-value pairs. There are an enormous number of objects in a large-size SAI. If relationship prediction is performed for all object pairs, more than $N(N-1)$ triplets will appear for $N$ objects in the case of one-to-many relationships (e.g., in the STAR dataset where the number of objects in one image is up to 6,800, the possible triplets will be more than 40,000,000), which will cause memory overflow under common computational resources. In fact, not all object pairs are of interest to us, and only a few of them contain high-value relationships. Therefore, it will be a new challenge to select high-value pairs from the numerous pairs for the SGG task in large-size VHR SAI.
@@ 3.2.2 Data Statistics and Visualization
The statistical results of object and relationship annotations from the STAR dataset are shown in Fig. 3(a) and Fig. 3(b), respectively. It can be seen that there is an obvious long-tail effect for both object and relationship categories in the STAR dataset, which is also consistent with reality. The highest-frequency objects are common small objects such as “cars”, “airplanes”, “tanks” and “boats”. The highest-frequency relationships are mainly related to the parking states, such as ”parked alongside with”, “parking in the different apron with”, “parallelly parked on” and “parking in the same apron with”.
@@ 4.1 Oriented Object Detection in Large-size VHR SAI
OBD in large-size SAI is an indispensable basis for SGG in large-size SAI, which directly affects the prediction accuracy of the triplets. Compared with small-size or low-resolution imagery, large-size VHR SAI can not only accommodate more complete large objects (e.g., runways, docks), but also greatly retain important details of small objects (such as the head and tail of a car, the head and tail of an airplane), which can help to mine more high-value knowledge in SAI. In processing large-size VHR SAI, mainstream deep learning-based OBD methods, whether supervised [53, 54, 55, 56] or weakly supervised [57, 58, 59, 60], adopt a cropping strategy [4, 61], which results in large objects, such as runways, being cut off. In addition to the cropping strategy, some OBD methods tackle the original large-size imagery with a fixed-window down-sampling strategy [62], resulting in the loss of massive detail information of the image. Recently, HBD-Net for single-class bridge detection in large-size VHR SAI is proposed [26], and the SDFF architecture in HBD-Net effectively solves the large scale variations in single-class bridge detection by fusing multi-scale contexts via the dynamic image pyramid (DIP) of the large-size image. However, this single-class OBD only considers the intra-class variations of the objects, which is not applicable in the multi-class OBD task with both intra-class and inter-class variations.
To address the aforementioned problem, it extends HBD-Net to the holistic multi-class OBD in large-size VHR SAI, named HOD-Net whose total loss of the HBB-based and OBB-based detectors in large-size VHR SAI is defined as:
$\mathcal{L}_{O}$ $=\sum_{m=1}^{\mathcal{M}}\left(\frac{1}{\Gamma^{m}}\!\sum_{i\in\Delta^{m}}\!\mathcal{L}_{i}^{cls}\!+\!\frac{1}{{\Gamma^{m}}^{+}}\sum_{j\in{\Delta^{m}}^{+}}\!w^{reg}_{j}\mathcal{L}_{j}^{reg}\right),$ (2)
$\mathcal{L}_{H}$ $=\sum_{m=1}^{\mathcal{M}}\left(\frac{1}{\Gamma^{m}}\!\sum_{i\in\Delta^{m}}\!\mathcal{L}_{i}^{cls}\!+\!\frac{1}{{\Gamma^{m}}^{+}}\!\sum_{j\in{\Delta^{m}}^{+}}\!\mathcal{L}_{j}^{reg}\right),$ (3)
@@ 4.2 Pair Proposal Generation via Adversarial Reconstruction
The training objective of $PED_{1}$ is to minimize the reconstruction error $\left\|X_{i}-X_{i}^{1}\right\|_{2}$ to learn the latent representation of the data, and the training objective of $PED_{2}$ is to reduce the reconstruction error $\left\|X_{i}-X_{i}^{2}\right\|_{2}$ to a minimum. As the iterations $n$ are constantly changing, the loss is:
@@
4.3 Relation Prediction with Context-Aware Messaging
@@ 4.3.1 Progressive Bi-context Augmentation Network
Figure 7: The proposed PBA module: the details of messaging and fusion for one progressive context augmentation of entity-relationship.
@@
where $\sigma(\cdot)$ denotes the sigmoid activation function [66]. $\alpha^{ee},\alpha^{rs}$ and $\alpha^{ro}$ represent the importance of messaging from entity to entity, relationship to subject, and relationship to object, respectively. By learning to adjust $\alpha$ [65], each iteration leads to a change in attention and affects subsequent iters. $w^{ee},w^{rs}$ and $w^{ro}$ denote the weight matrix of the directed messaging from entity to entity, relationship to subject, and relationship to object, respectively. ${F}_{i}^{e}$ and ${F}_{i}^{r}$ are entity and relationship representation, respectively.
@@
Global-Local Feature Fusion. Introducing reliable global information helps to better recognize relationships, but more fine-grained discriminative features mainly come from local. Therefore, we adopt a global-local feature fusion strategy to capture reliable global information and maintain the local more discriminative features. Specifically, the context augmented feature of entity ${F}_{i+1}^{e}$ can be obtained by fusing the context-aware feature of entity $\widehat{F}_{i+1}^{e}$ from global and the entity feature ${F}_{0}^{e}$ from local. Similarly, the context augmentation feature of relationship ${F}_{i+1}^{r}$ can be obtained. After $L$ iters, the predicted relationship representations $Rel$ with context-aware augmentation are updated. Note that the predicted relationship representations in Sec. 4.3.1 are optimized and updated simultaneously with the prototype representations in Sec. 4.3.2.
@@ 4.3.2 Prototype-guided Relationship Learning
Prototype learning allows SGG models to focus on all categories equally in training, and is effective in alleviating the category imbalance problem of datasets [45]. However, previous linear classifier-based methods often lack the ability to model data at different semantic levels and suffer from a lag in representational capabilities (e.g., $<$ship, away from, dock$>$ and $<$ship, approach, dock$>$ correspond to different relationship labels, but it is clear that $<$ship, away from, dock$>$ and $<$ship, approach, dock$>$ are closer than $<$truck, away from, gas station$>$ in the visual representation). We propose a prototype-guided relationship learning method that utilizes category labels as prompts to help generate learnable relationship prototypes, and prototype representations updated dynamically by instance-level constraints and prototype-level constraints to help the SGG model enhance the discriminability between prototypes.
@@
$\mathcal{L}_{IC}$ $=-\log\frac{\exp(<\overline{r},\overline{p}>/\tau)}{\textstyle\sum_{i=0}^{C}\exp(<\overline{r},\overline{p}_{i}>/\tau)},$ (8)
$\mathcal{L}_{ID}$ $=\max(0,q^{+}-q^{-}+\gamma_{1}),$ (9)
@@
The overall training loss $\mathcal{L}$ of our RPCM is as follows, without any bias on the weight of each term:
@@
In the inference stage, the prototype type with the highest cosine similarity $s_{i}$ is used for relationship prediction:
@@
TABLE II: Baseline results (%) of HBB-based and OBB-based detectors on STAR test set. b_b denotes boarding_bridge, l_t denotes lattice_tower, s_l is ship_lock, and g_d represents gravity_dam. All experiments are based on the standard ‘1x’ (12 epochs) training schedule. means that Swin-L [67] is used and others indicates ResNet50 [68] as the backbone. Underline indicates the base detector for subsequent lines.
@@
TABLE III: Results (%) of different training strategies on STAR test set. b_b denotes boarding_bridge, l_t denotes lattice_tower, s_l is ship_lock, and g_d represents gravity_dam. All experiments are based on the standard ‘1x’ (12 epochs) training schedule. Swin-L [67] is adopted as the backbone.
@@
Figure 8: The visualization results of OBB and HBB detection on the STAR dataset using the HOD-Net and comparison OBD methods.